OpenStudy (anonymous):

Can someone help me? I've been stuck on this problem and a few others for a long time. Solve the system of equations by substitution. What is the solution for x? (1 point) 2x + y = 1 4x + 2y = -2 x = 0 x = 2 There is no x value, as there is no solution. x can be any value as, there is an infinite number of solutions.

OpenStudy (danjs):

2x + y = 1 4x + 2y = -2 Substitution means, solve one of the equations for a variable, like y = ..., then put that in the second equation

OpenStudy (danjs):

For example, THe first equation : 2x + y = 1, solving for y y = -2x + 1

OpenStudy (anonymous):

Well, math is one of my really bad subjects. So I don't even know where to start to try and figure this out.

OpenStudy (danjs):

No Solution - The two equations of lines, are parallel lines, they dont cross Infinite Solutions - They are the same line overlapping each other

OpenStudy (danjs):

Alright, no problem, we can go through it a step at a time

OpenStudy (anonymous):

Okay thank you

OpenStudy (danjs):

2x + y = 1 4x + 2y = -2 Those are the original equations in the problem, When they ask you about solutions to these equations, they want to know, where do they cross each other on the graph , the x-y plane...

OpenStudy (danjs):

That happens at a point in the plane (x,y) x= something and y= something

OpenStudy (danjs):



OpenStudy (danjs):

That is what these types of problems can look like, they want you to find that solution point (x,y) if there is one

OpenStudy (danjs):

Just where 2 lines cross each other, if they do

OpenStudy (anonymous):

This is confusing.. but I am trying to follow along.

OpenStudy (danjs):

I am just describing what these types of problems look like on a graph. Each equation is just a line. The solution to them x= something and y= something is where the 2 lines cross each other.

OpenStudy (anonymous):

I have four more questions basically just like this one... so I am trying to understand it so I can answer the other ones. So far I am not getting anywhere. But thank you for trying to explain

OpenStudy (danjs):

2x + y = 1 4x + 2y = -2 Here is the mechanical way to solve them, without knowing anything , just the process steps for substitution.....

OpenStudy (anonymous):

Except the others ones aren't asking to substitute, they're just asking to solve the following system of equations.

OpenStudy (danjs):

Solve one equation for y = First equation 2x + y = 1 solve for y y = -2x + 1

OpenStudy (danjs):

Good so far?

OpenStudy (anonymous):

Yes, but with the other equation do I just do Y= 4x + 2 + -2.. 0r? becuase the 2 is next to the Y

OpenStudy (anonymous):

Sorry If I am asking dumb questions

OpenStudy (danjs):

Now that you solved the first equation for y, y = -2x + 1 Substitute that y into the second equation wherever you see a 'y' 4x + 2y = -2 4x + 2*(-2x+1) = -2 Now you have the second equation , with just an X variable and no Y 4x + 2*(-2x+1) = -2

OpenStudy (danjs):

Now that you substituted , and have an equation with only an X variable, you can solve it for X = .... 4x + 2(-2x+1) = -2 4x -4x + 2 = -2 0x + 2 = -2 2 = -2 FALSE, this system of 2 equations, has ZERO solutions

OpenStudy (anonymous):

Okay, when you go through the work like that it starts to make more sense. But still there is a lot for me to look at to try and get where you find the answer. But it is making more sense so thank you

OpenStudy (danjs):

If you solve and find A false Statement, like 23 = 10 - there are no solutions A true statement, like 15 = 15 - there are infinite many solutions A value for x, then you have to find y, and the solution is the pair (x,y)

OpenStudy (anonymous):

So would the answer just be x=0? or is there still more work we have to do

OpenStudy (danjs):

Ok, that is why i wanted to show you the graph, Each of the equations describes a LINE on the graph, If there are no solutions, they are parallel lines, they do not cross each other If there are Infinite solutions , they are the same line, overlapping each other IF there is ONE solution, It is the point they cross on the graph

OpenStudy (danjs):

For this one we solved and just got a false statement 2 = -2 False --- that means there are no solutions for x

OpenStudy (danjs):

If you graphed both of the lines, they would be parallel and not cross each other

OpenStudy (anonymous):

So there's just no solution at all. Okay..

OpenStudy (danjs):

right

OpenStudy (anonymous):

Thank you. I am starting to understand a little better now.. Do you think you could help me with a few more please?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

sure, post up another thread, and put my name in it